I don't know how to solve a quadratic equation by graphing. I know how by using the quad. equation. this is the problem I need to solve by graphing. 5x^2 + 20x+15=0
Answer by checkley79(3341)

5x^2 + 20x+15=0
5(X^2+4X+3)=0
5(X+3)(X+1)=0
X+3=0
X=-3 ANS.
X+1=0
X=-1 ANS.
PROOFS:
5*-3^2+20*-3+15=0
5*9-60+15=0
45-60+15=0
60-60=0
0=0
5*-1^2+20*-1+15=0
5-20+15=0
20-20=0
0=0